OpenStudy (anonymous):

Find the derivative of the function. f(x)=6x^3-2x-5

OpenStudy (unklerhaukus):

\[f(x)=6x^3-2x-5\] \[\frac{\text df(x)}{\text dx}=18x^2-2\]

OpenStudy (anonymous):

why 18? why not 12?

OpenStudy (anonymous):

3*6 = 18

OpenStudy (anonymous):

oh so it's the exponent times the coefficient..

OpenStudy (anonymous):

yes and take away one from the exponent after

OpenStudy (unklerhaukus):

\[\frac {\text d }{\text dx}\left(x^n\right)=nx^{n-1}\]

OpenStudy (anonymous):

isn't there supposed to be a +C to replace the 5? or is that antiderivatives?

OpenStudy (anonymous):

thats anti

OpenStudy (anonymous):

that's**

OpenStudy (anonymous):

oh all right. Thanks guys.

OpenStudy (anonymous):

the is for when you don't know what the constant is but when you take a derivative you know what the constant is

OpenStudy (unklerhaukus):

c is constant of integration,

OpenStudy (anonymous):

the 'c'

OpenStudy (anonymous):

oh i see.

OpenStudy (anonymous):

Know the product formula?

OpenStudy (anonymous):

yea what they mentioned above is the shortcut i was confused about. They cleared it up for me. Thanks anyway =D

OpenStudy (anonymous):

No problem, that is the product formula.